# Version dated: November 6, 2025

### v1

@@ 5 Discussion
A critical aspect of our methodology is the use of an efficient inferential scheme based on HMC with surrogate trajectories. Given the substantial computational burden associated with exact gradient computations for matrix exponentials in partially observed CTMCs, we leverage an approximation of the matrix exponential derivative (didier_surprising_2024). This approximation dramatically reduces the computational complexity of the log-posterior gradient from ${\cal O}(K^{5})$ to ${\cal O}(K^{2})$ with respect to all log-rates simultaneously, making inference tractable for larger state spaces. The theoretical underpinning provided by li_neural_2019 ensures that, with a Metropolis-Hastings correction step, the exact posterior remains the stationary distribution of the Markov chain despite using an approximate gradient. Beyond methodological advances, we contribute directly to the field of Bayesian phylogenetics by integrating the proposed methods into the widely used Beast X software. This implementation is intended to facilitate broader adoption within the scientific community and promote the exploration of covariate effects in phylogeographic models.
@@
Future work could extend the framework to account for uncertainty in the covariates themselves, which are currently treated as fixed and known. Incorporating covariate uncertainty may yield more robust inference, particularly in settings where the external information is noisy or incomplete. Additionally, while the current approximation significantly enhances computational efficiency, further research could explore alternative scalable gradient approximations or inference strategies to support even larger or more complex state spaces.
In summary, the proposed framework offers a robust and scalable solution for Bayesian inference of CTMCs. By providing a principled way to model nonlinear covariate effects and employing computationally efficient gradient approximations, this work extends the capabilities of current models, allowing for deeper insights into the dynamics of stochastic processes.
@@ Acknowledgments
6 Appendix
6.1 Computing the tree-data likelihood gradient
@@
To compute the gradient of the tree data likelihood with respect to the normalized rates, we leverage the chain rule to single out the contributions to the gradient of the rates specific to each branch and the $\boldsymbol{\pi}_{\text{root}}$. To accomplish this, we introduce a set of auxiliary branch- and rate-specific functions: $\eta_{uij}(\lambda_{ij})=\lambda_{ij}$. Since $\boldsymbol{\pi}_{\text{root}}$ also possibly depends on all the rates, we extend the definition of the functions above such as $\eta_{2N-1,ij}(\lambda_{ij})=\eta_{\text{root},ij}(\lambda_{ij})=\lambda_{ij}$, where 2N12N-1 is the root node.
Therefore, substituting each rate at each branch and at the root with the corresponding auxiliary functions, we can apply the chain rule as follows:
$\frac{\partial}{\partial\lambda_{ij}}=\sum_{u=1}^{2N-1}\frac{\partial}{\partial\eta_{uij}}\frac{\partial\eta_{uij}}{\partial\lambda_{ij}}=\sum_{u=1}^{2N-1}\frac{\partial}{\partial\eta_{uij}},$ (31)
@@
where the second equality stems from $\frac{\partial\eta_{uij}}{\partial\lambda_{ij}}=1$. Using this fact as well as the characterization of the tree data likelihood based on the pre- and post-order partial likelihood shown in Equation 25, and some further mathematical manipulations, we obtain
@@
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\frac{\partial}{\partial\eta_{uij}}{\mathbf{P}}\hskip-1.99997pt\left(t_{u}\right)^{\prime}\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(t_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\frac{\partial}{\partial\eta_{uij}}{\mathbf{P}}\hskip-1.99997pt\left(t_{u}\right)^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\eta_{uij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\sum_{ab}\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\lambda_{ab}}\frac{\partial\lambda_{ab}}{\partial\eta_{uij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}},$ (32)
where the second line follows since $\mathbf{p}_{u}$ does not depend on the branch connecting node $u$ to its parent and by applying the pre-order pruning algorithm formula (27). Moreover, $\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$ represents the contribution to the derivative at the root node. In the third line, we define for notational convenience
$\tilde{\mathbf{q}}_{\text{pa}(u)}:=\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(t_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]$ (33)
@@
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{t_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}.$ (34)

### v2

@@ 5 Discussion
A critical aspect of our methodology is the use of an efficient inferential scheme based on HMC with surrogate trajectories. Given the substantial computational burden associated with exact gradient computations for matrix exponentials in partially observed CTMCs, we leverage an approximation of the matrix exponential derivative (Didier et al., 2024). This approximation dramatically reduces the computational complexity of the log-posterior gradient from ${\cal O}(K^{5})$ to ${\cal O}(K^{2})$ with respect to all log-rates simultaneously, making inference tractable for larger state spaces. The theoretical underpinning provided by Li et al. (2019) ensures that, with a Metropolis-Hastings correction step, the exact posterior remains the stationary distribution of the Markov chain despite using an approximate gradient. Beyond methodological advances, we demonstrate the utility of the proposed framework in Bayesian phylogenetics by integrating the proposed methods into the widely used Beast X software. This implementation supports reproducible application of the framework and enables the exploration of covariate effects in phylogeographic models.
@@
Future work could extend the proposed framework to account for uncertainty in the covariates, which are currently treated as fixed and known. This could be achieved through hierarchical or measurement-error models that explicitly propagate covariate uncertainty into inference on the infinitesimal rates, potentially improving robustness in settings where external information is noisy or incomplete. In addition, while the current gradient approximation substantially reduces computational cost, further research could investigate alternative scalable gradient approximations or complementary inference strategies to accommodate even larger state spaces or more complex dependency structures.
In summary, this work introduces a flexible and scalable Bayesian framework for inference on continuous-time Markov chains with covariate-dependent rates. By enabling nonlinear covariate effects and supporting efficient gradient-based computation, the proposed methodology broadens the class of CTMC models that can be practically analyzed and provides a foundation for further methodological development.
@@ Acknowledgments
SUPPLEMENT
S1 Gradient with respect to the auxiliary parameters in partially observed CTMCs
Let $\mathbf{Q}=(q_{ij})_{ij}$ be the infinitesimal matrix of a CTMC from which we sample $N$ observations $\mathbf{y}=(y_{1},y_{2},\dots,y_{N})$. We introduce real-valued auxiliary variables $\theta_{ij}=g^{-1}(q_{ij})$ where $g$ is a non-negative differentiable one-to-one function. When a CTMC is partially observed, to ensure identifiability, we adopt the normalization constant $\psi=\sum_{i\neq j}\pi_{i}q_{ij}$, where $\boldsymbol{\pi}$ is a distribution vector over the CTMC states. Then, we define a normalized infinitesimal generator $\mathbf{\Lambda}=(\lambda_{ij})_{ij}$ element-wise by $\lambda_{ij}=\frac{q_{ij}}{\psi}$. $\mathbf{\Lambda}$ characterizes a new CTMC in which time and infinitesimal rates are identifiable. In this new normalized formulation, the likelihood $p\left(\mathbf{y}\right)$ depends directly only on the normalized rates $\lambda_{ij}$. Therefore, differentiating $p\left(\mathbf{y}\right)$ with respect to the auxiliary parameters $\theta_{ij}$ needs to account for their effect on each of the normalized rates. We now formally derive the likelihood gradient and evaluate it when the link function $g$ is exponential.
First, we need to apply the chain rule on $p\left(\mathbf{y}\right)$, separating the effect of $\theta_{ij}$ on each $\lambda_{ij}$, that is:
$=\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\frac{1}{\psi}\frac{\partial q_{uv}}{\partial\theta_{ij}}-\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}q_{uv}\frac{1}{\psi^{2}}\frac{\partial\psi}{\partial\theta_{ij}}.$ (S1.1)
Noticing that $\frac{\partial q_{uv}}{\partial\theta_{ij}}=0$ for all $uv\neq ij\text{ and }ii$, we can re-write the first summation in Equation S1.1 as:
$\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\frac{1}{\psi}\frac{\partial q_{uv}}{\partial\theta_{ij}}$ $=\left(\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}-\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ii}}\right)\frac{1}{\psi}\frac{\partial q_{ij}}{\partial\theta_{ij}}=\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}\frac{1}{\psi}\frac{\partial q_{ij}}{\partial\theta_{ij}}$ (S1.2)
where $\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}:=\left(\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}-\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ii}}\right)$ is the difference between the partial derivative with respect to a non-diagonal element and its corresponding diagonal element. We can then insert Equation S1.2 into Equation S1.1 and input the exponential link function $q_{ij}=e^{\theta_{ij}}$ such that:
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\theta_{ij}}$ $=\left[\Delta\hskip-2.5pt\left(\mathbf{y}\right)_{ij}-\left(\sum_{uv}\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{uv}}\lambda_{uv}\right)\pi_{i}\right]\lambda_{ij},$ (S1.3)
recalling that $\frac{\partial\psi}{\partial\theta_{ij}}=\pi_{i}q_{ij}=\pi_{i}\psi\lambda_{ij}$. Notice that these expressions only depend on the distribution vector $\boldsymbol{\pi}$, the normalized rates $\lambda_{ij}$, and the gradient of $p\left(\mathbf{y}\right)$ with respect to each normalized rate. We can therefore evaluate the gradient directly using the matrix exponential approximation proposed in the main text.
S2 A Comparison with automatic differentiation
Automatic differentiation (AD) methods have become a cornerstone of modern scientific computing, particularly in fields like machine learning and optimization. By providing efficient and exact ways to compute derivatives, they have enabled the development of complex models that were previously intractable (Margossian, 2019). In general, AD algorithms are divided into two main modes: forward mode and reverse mode. Forward-mode AD is most efficient when the function of interest depends on a small number of inputs but produces many outputs, whereas reverse-mode AD is preferable when there are many input variables and few output quantities (Baydin et al., 2018). Given our framework’s structure—with many input variables (the rates) and a single scalar output (the joint distribution)—reverse-mode AD constitutes the natural choice for gradient evaluation. This method requires a forward pass to compute the function value, which scales as ${\cal O}(K^{3})$, followed by a single backward pass to obtain all parameter gradients simultaneously. The backward pass requires computing the adjoint of the derivative of the matrix exponential, an operation of comparable complexity, also ${\cal O}(K^{3})$.
Following the same reasoning as in the previous section, within the context of CTMCs, evaluating the likelihood gradient requires at least N1N-1 adjoint computations, resulting in an overall computational complexity of ${\cal O}(NK^{3})$. Although this represents a substantial improvement over performing a full sequence of exact gradient evaluations for each individual rate parameter, it remains an order of magnitude more expensive than directly applying the proposed approximate gradient. The gap in efficiency becomes even more pronounced when considering the large $K$-invariant constant factor embedded in the ${\cal O}(NK^{3})$ cost. Specifically, computing the adjoint using the exact gradient from Equation 12 of the main text requires the evaluation of a matrix exponential of dimension $2K\times 2K$, whose cubic scaling implies a computational burden proportional to $(2K)^{3}=8K^{3}$. Moreover, even the most efficient algorithms for computing a matrix exponential require work equivalent to about $29$ matrix-matrix multiplications (Najfeld and Havel, 1995).
S3 Computing the tree-data likelihood gradient
S3.1 Background
Let $\mathbf{y}$ be data observed at tips of an $N$-taxa tree and $v$ be an internal node. Let $y_{\lfloor v\rfloor}$ collect all descendants of $v$ and $y_{\lceil v\rceil}$ collect the non-descendants. Leveraging conditional independence, the tree data likelihood can be computed via:
$p\left(\mathbf{y}\right)={p}\hskip-1.00006pt\left(y_{1},\dots,y_{N}\right)$ $=\sum_{y_{v}\in\mathbf{S}}{p}\hskip-1.00006pt\left(y_{\lfloor v\rfloor},y_{v}\right){p}\hskip-1.00006pt\left(y_{\lceil v\rceil}\hskip-1.99997pt\mid\hskip-1.00006pty_{v}\right)=\mathbf{p}_{v}^{{}^{\prime}}\mathbf{q}_{v},$ (S3.1)
where $\mathbf{p}_{v}$ is the post-order partial likelihood vector containing ${p}\hskip-1.00006pt\left(y_{\lfloor v\rfloor}\hskip-1.99997pt\mid\hskip-1.00006pty_{v}=s_{k}\right)$ for each state $s_{k}\in\mathbf{S}$, and $\mathbf{q}_{v}$ is the pre-order partial likelihood vector containing ${p}\hskip-1.00006pt\left(y_{\lceil v\rceil},y_{v}=s_{k}\right)$ for each state.
Let $u_{1}$ and $u_{2}$ be the two children of node $v$ with $b_{u_{1}}$ and $b_{u_{2}}$ representing the branches connecting them to the parent. Then, the post-order partial likelihood can be derived iteratively from the tips to the root using
$\mathbf{p}_{v}$ $={\mathbf{P}}\hskip-1.99997pt\left(b_{u_{1}}\right)\mathbf{p}_{u_{1}}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{u_{2}}\right)\mathbf{p}_{u_{2}},$ (S3.2)
where $\circ$ represents the Hadamard or element-wise product. Besides, the algorithm for the pre-order partial likelihoods runs from the root to the tips applying
$\mathbf{q}_{u_{1}}$ $={\mathbf{P}}\hskip-1.99997pt\left(b_{u_{1}}\right)^{\prime}\left(\mathbf{q}_{v}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{u_{2}}\right)\mathbf{p}_{u_{2}}\right).$ (S3.3)
S3.2 Gradient Computations
@@
To compute the gradient of the tree data likelihood with respect to the normalized rates, we leverage the chain rule to single out the contributions to the gradient of the rates specific to each branch and the distribution at the root $\boldsymbol{\pi}_{\text{root}}$. To accomplish this, we introduce a set of auxiliary branch- and rate-specific functions: $\eta_{uij}(\lambda_{ij})=\lambda_{ij}$. Since $\boldsymbol{\pi}_{\text{root}}$ also possibly depends on all the rates, we extend the definition of the functions above such that $\eta_{2N-1,ij}(\lambda_{ij})=\eta_{\text{root},ij}(\lambda_{ij})=\lambda_{ij}$, where 2N12N-1 is the root node. Therefore, substituting each rate at each branch and at the root with the corresponding auxiliary functions, we can apply the chain rule as follows:
$\frac{\partial}{\partial\lambda_{ij}}=\sum_{u=1}^{2N-1}\frac{\partial}{\partial\eta_{uij}}\frac{\partial\eta_{uij}}{\partial\lambda_{ij}}=\sum_{u=1}^{2N-1}\frac{\partial}{\partial\eta_{uij}},$ (S3.4)
where the second equality stems from $\frac{\partial\eta_{uij}}{\partial\lambda_{ij}}=1$.
@@
Using this fact as well as the characterization of the tree data likelihood based on the pre- and post-order partial likelihood shown in Equation S3.1, and some further mathematical manipulations, we obtain
@@
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\frac{\partial}{\partial\eta_{uij}}{\mathbf{P}}\hskip-1.99997pt\left(b_{u}\right)^{\prime}\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\frac{\partial}{\partial\eta_{uij}}{\mathbf{P}}\hskip-1.99997pt\left(b_{u}\right)^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\eta_{uij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$
$=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\sum_{ab}\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\lambda_{ab}}\frac{\partial\lambda_{ab}}{\partial\eta_{uij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}},$ (S3.5)
where the second line follows since $\mathbf{p}_{u}$ does not depend on the branch connecting node $u$ to its parent and by applying the pre-order pruning algorithm formula (Equation S3.3). Moreover, $\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}$ represents the contribution to the derivative at the root node. In the third line, we define for notational convenience
$\tilde{\mathbf{q}}_{\text{pa}(u)}:=\left[\mathbf{q}_{\text{pa}(u)}\circ{\mathbf{P}}\hskip-1.99997pt\left(b_{\text{sibl}(u)}\right)\mathbf{p}_{\text{sibl}(u)}\right]$ (S3.6)
@@
$\frac{\partial p\left(\mathbf{y}\right)}{\partial\lambda_{ij}}$ $=\sum_{u=1}^{2N-2}\mathbf{p}_{u}^{\prime}\left[\frac{\partial e^{b_{u}\mathbf{\Lambda}}}{\partial\lambda_{ij}}\right]^{\prime}\tilde{\mathbf{q}}_{\text{pa}(u)}+\mathbf{p}_{2N-1}^{\prime}\frac{\partial}{\partial\lambda_{ij}}\boldsymbol{\pi}_{\text{root}}.$ (S3.7)